OpenStudy (anonymous):

when the hypotenuse is 11.1 and one leg of the right triangle is 7.2 whats the other leg?

OpenStudy (amistre64):

what does the pythag thrm give us?

OpenStudy (anonymous):

Okay so. The pythagorean theorem tells us that \[a^2+b^2=c^2\].  Let's substitute the numbers. Let: a=7.2 b=x c=11.1 Now substitute these numbers into the equation! \[7.2^2+x^2=11.1^2\] \[x^2=11.1^2-7.2^2\] Now just solve that equation! :)